In the laboratory you dissolve 12.6 g of iron(II) nitrate in a volumetric flask and add water to a total volume of 500 mL.

1. What is the molarity of the solution? ______ M.
2. What is the concentration of the iron(II) cation? ______M.
3. What is the concentration of the nitrate anion? _______ M.
1
Expert's answer
2019-03-15T04:19:59-0400

Fe(NO3)2 -> Fe2+ + 2NO3-

n=m/M

n(Fe(NO3)2) = 12.6 g/179.87 g/mol = 0.07 mol

  1. c(Fe(NO3)2) = 0.07 mol/0.5 L = 0.14 mol/L = 0.14 M
  2. According to equation n(Fe2+) = n(Fe(NO3)2)= 0.07 mol

c(Fe2+) = 0.07 mol/0.5 L = 0.14 mol/L = 0.14 M

3.According to equation mole ratio n(Fe(NO3)2):n(NO3-) = 1:2, then n(NO3-) = 2*n(Fe(NO3)2) = 2*0.07 mol = 0.14 mol

c(NO3-) = 0.14 mol/0.5L = 0.28 mol/L = 0.28M